how do i work out a formula when asked the question "i played 34 basketball games and won 8 more than i lost,...how many games did i win
thank you

Answer by stanbon(75887)

work out a formula when asked the question "i played 34 basketball games and won 8 more than i lost,...how many games did i win
----------------
Let the number of your wins be "x"
Then the number of your losses is "x+8".
-----------------------
EQUATION:
wins + losses = 34
x + (x+8) = 34
2x +8 = 34
2x = 26
x = 13 (number of wins)
x+8 = 21 (number of losses)
===============================
Cheers,
Stan H.